# The quenched structured coalescent for diploid population models on finite graphs with large migrations and uneven offspring distributions

### v1

@@ Abstract
In this work we describe a new model for the evolution of a diploid structured population backwards in time that allows for large migrations and uneven offspring distributions. The model generalizes both the mean-field model of Birkner et al. [Electron. J. Probab. 23: 1-44 (2018)] and the haploid structured model of Möhle [Theor. Popul. Biol. 2024 Apr:156:103-116]. We show convergence, with mild conditions on the joint distribution of offspring frequencies and migrations, of gene genealogies conditional on the pedigree to a time-inhomogeneous coalescent process driven by a Poisson point process $\Psi$ that records the timing and scale of large migrations and uneven offspring distributions. This quenched scaling limit demonstrates a significant difference in the predictions of the classical annealed theory of structured coalescent processes. In particular, the annealed and quenched scaling limits coincide if and only if these large migrations and uneven offspring distributions are absent. The proof proceeds by the method of moments and utilizes coupling techniques from the theory of random walks in random environments. Several examples are given and their quenched scaling limits established.
@@ 1 Introduction
Tyukin (2015) established the first result for quenched coalescents, showing that in a well-mixed population in the absence of uneven offspring distributions, gene genealogies of unlinked loci really are independent in the large population limit. However, the works Diamantidis et al. (2024); Alberti et al. (2025) then showed the effects of uneven offspring distributions in the absence of selfing on gene genealogies. The pedigree leaves a persistent imprint on the dynamics of these gene genealogies through prodigious progenitors, individuals whose genetic contributions are shared across the genome. The limiting quenched genealogy becomes time inhomogeneous, evolving with an ordinary Kingman-like background punctuated by rare generations in which prodigious individuals contribute a non-vanishing proportion of the genetic material of the population even in the infinite population limit. Under the same assumptions of Birkner et al. (2018) that yield an annealed multiple merger coalescent, the quenched coalescent converges to a process that is driven by a Poisson point process recording the timing and structure of exceptional reproductive generations of the pedigree. Between exceptional generations, lineages merge only in binary mergers at a constant rate. At exceptional generations, many lineages may merge simultaneously, with the merger pattern determined by the ordered offspring frequencies. This identifies the precise sense in which widely separated loci are coupled; they share the same record of macrosopic pedigree events, these prodigious progenitors. This Poisson driven viewpoint is closely related to the limiting object in the present paper.
@@
Much like the well-mixed case, the standard structured coalescent describes an averaged single locus genealogy. It does not condition on the single realized history of the pedigree. In structured populations this distinction becomes especially natural. Large migration waves can occur on the same time scale as coalescence and can affect many lineages at once. Such events are recorded in the pedigree and are shared across loci. They therefore create an additional source of time inhomogeneity and cross locus dependence that is invisible to purely annealed descriptions. Wilton et al. (2017) provided simulation results that demonstrated the importance of the pedigree in structured populations, where deviations from standard coalescent predictions can persist longer than in well-mixed populations due to the specific history of migration events captured by the pedigree.
@@ 1.1 The contributions of this paper
We control these joint laws by constructing a coupling of the $l$ conditionally independent copies on the same pedigree, in the same spirit as Alberti et al. (2025), who built on earlier work of Bolthausen and Sznitman (2002); Birkner et al. (2012). The coupling makes transparent the two mechanisms that drive the limit. Firstly, genuinely large pedigree events that change the coalescent state, such as macroscopic migration waves or unusually large families, are rare on the coalescent time scale. However, they affect all copies in nearly the same way. In particular, the different copies jump to corresponding new states with asymptotically matching probabilities. Second, during the overwhelming majority of small generations, the copies experience the “neutral” migration patterns described above, and their jump decisions become asymptotically independent across copies. This approximate independence follows from straightforward moment bounds. These two observations form the basis for a “separation of scales” argument in which we bridge these two regimes to demonstrate the desired convergence of the joint law. We note here that there is an implicit separation of scales argument in the $\varepsilon$-naive-coalescent of Alberti et al. (2025), though the total form of the argument is distinct.
@@ 2.2 The pedigree and its ancestral process
Let $F$ be a topological space. A partition with type in $F$ is a pair $(\xi,f)$, where $\xi\in\mathcal{E}_{n}$ and
@@
  • the blocks containing $i$ and $j$ are paired in $\xi$ if and only if $\widehat{X}_{i}^{N}(k)=\widehat{X}_{j}^{N}(k)$,
@@ 3 An annealed scaling limit
for every $(v,w)$ in $E$.
@@
Blocks are identified if their points lie in the same deme and in the same interval determined by $x^{v}$. Denote by $q_{n}(x,m)(\xi,\eta)$ the probability that an element $\xi$ coalesces into the element $\eta$, conditional on $(x,m)$.
@@ 4 Main results
The proof essentially proceeds by the method of moments. This is possible, due to a corollary of Fan et al. (2025), which guarantees that quenched convergence in finite-dimensional distribution of càdlàg paths is enough to guarantee weak convergence in distribution of quenched laws when the space the paths take values in are compact. We make this reasoning precise in the following lemma.
@@
$\mathbb{E}\left[\mathbb{P}\left(\overline{\chi}^{N,n}\in C(\vec{t},\vec{\xi})\,\mid\,\mathcal{A}_{N}\right)^{l}\right]=\mathbb{P}\left(\left(cd_{V}\left(\overline{\chi}_{i}^{N,n}\right)\right)_{i=1}^{l}\in C(\vec{t},\vec{\xi})^{l}\right)$
@@
At the same time, when we restrict ourselves to the rare jumps where $(\widetilde{\mathcal{V}}(k),m(k))\in B(\varepsilon)^{c}$ we have that the transitions of the conditionally independent realizations are really asymptotically independent.
@@
Suppose, as $N\to\infty$, that Assumptions 3.2, 3.3, 3.4, 3.7, and 3.12 hold. Let $P_{N,l}(t)$ denote the transition kernel of $\left(cd_{V}\left(\overline{\chi}_{i}^{N,n}\right)\right)_{i=1}^{l}$ at time $t$. Then, for any fixed $T>0$, there is a continuous function $g_{T}:\mathbb{R}_{+}\to\mathbb{R}_{+}$ with $g(0)=0$ such that
@@
Finally, we characterize the annealed joint law of $l$ realizations $\left(\chi_{i}^{n}\right)_{i=1}^{l}$ of a $\Psi$-driven $n$-$(\kappa,\mu)$-coalescent $\chi^{n}$ that are conditionally independent with respect to $\Psi$. We show that $\left(\chi_{i}^{n}\right)_{i=1}^{l}$ has infinitesimal generator $\mathcal{L}$
@@ 6.1 A two-deme Wright-Fisher model with neutral migration
$m_{(1,2)}=m_{(2,1)}\sim\frac{1}{N}\text{\emph{Binomial}}\left(N;\frac{\sigma}{2}\right)$
@@ 6.2 A two-deme modified Wright-Fisher model with large offspring numbers
$m_{(1,2)}=m_{(2,1)}\sim\frac{1}{N_{\gamma}}\text{\emph{Binomial}}\left(N;\frac{N_{\gamma}}{N}\frac{\sigma}{2}\right)$
@@
$\kappa_{\gamma}:=\begin{cases}\left(1,1\right),&\gamma>1,\\ \left(\dfrac{1}{1+\phi\psi^{2}},\dfrac{1}{1+\phi\psi^{2}}\right),&\gamma=1,\\ \left(0,0\right),&0<\gamma<1,\end{cases}\qquad\mu_{\gamma,(1,2)}=\mu_{\gamma,(2,1)}:=\begin{cases}\mu,&\gamma>1,\\ \dfrac{\mu}{1+\phi\psi^{2}},&\gamma=1,\\ 0,&0<\gamma<1.\end{cases}$
@@
By a direct computation one obtains $\mu_{\gamma,(1,2)}=\mu_{\gamma,(2,1)}$ as stated in the proposition. Hence Assumption 3.7 holds.
It remains to identify the vague limit in Assumption 3.4. In a small generation, $\widetilde{\mathcal{V}}^{v}$ is Wright-Fisher reproduction, so its contribution vanishes on $\Delta^{V}\times[0,1]^{E}\setminus\{\mathbf{0}_{V,E}\}$ by (Birkner et al., 2018, Proposition 2.1). In a large generation in deme $v$, the $\lfloor\psi N\rfloor$ offspring of the prolific couple inherit, at a fixed locus, one of the two parental copies from each of the two parents independently; thus the four parental copies each contribute $\mathrm{Bin}\left(\lfloor\psi N\rfloor,1/2\right)$ gametes, and after normalisation by $2N$ this yields the limit $x_{\psi}$ in $\Delta$. Since large generations occur with probability $\varepsilon_{N}$ independently in the two demes, the event of simultaneous large generations has probability $\varepsilon_{N}^{2}=o(c_{N})$, and does not contribute in the $1/c_{N}$ scaling. Consequently,
@@
given by (6.3). This is exactly Assumption 3.4 with $\Phi=\Phi_{\gamma}$.
@@ 6.3 A two-deme model with random individual fitness and large migrations
$\mathbb{P}\left(W\geq z\right)\sim c_{W}z^{-\alpha}\qquad\text{as }z\to\infty$
@@
$\Lambda_{\alpha}(dy)=\frac{1}{\Gamma(2-\alpha)\Gamma(\alpha)}\,y^{1-\alpha}(1-y)^{\alpha-1}\,dy,\qquad\nu_{\alpha}(dy):=\frac{\Lambda_{\alpha}(dy)}{y^{2}}.$
@@
$\Phi^{\mathrm{fit}}:=c_{\alpha}\int_{(0,1)}\left(\delta_{\left(x_{y}^{(1)},0_{E}\right)}+\delta_{\left(x_{y}^{(2)},0_{E}\right)}\right)\,\nu_{\alpha}(dy),$
@@
by (Birkner et al., 2018, Lemma 2.2). In particular, Assumptions 3.2 and 3.3 hold. Moreover, since $c_{N}\gg 1/N$ for $\alpha\in(1,2)$, the Kingman contribution disappears under the $1/c_{N}$ time-rescaling, hence $\kappa=\left(0,0\right)$.

### v2

@@ Abstract
In this work we describe a new model for the evolution of a diploid structured population backwards in time that allows for large migrations and uneven offspring distributions, and that generalizes both the mean-field model of Birkner et al. [Electron. J. Probab. 23: 1-44 (2018)] and the haploid structured model of Möhle [Theor. Popul. Biol. 2024 Apr:156:103-116]. We show convergence, with mild conditions on the joint distribution of offspring frequencies and migrations, of gene genealogies conditional on the pedigree to a time-inhomogeneous coalescent process driven by a Poisson point process $\Psi$ that records the timing and scale of large migrations and uneven offspring distributions. This quenched scaling limit demonstrates a significant difference in the predictions of the classical annealed theory of structured coalescent processes. In particular, the annealed and quenched scaling limits coincide if and only if these large migrations and uneven offspring distributions are absent. The proof proceeds by the method of moments and utilizes coupling techniques from the theory of random walks in random environments. Several examples are given and their quenched scaling limits established.
@@ 1 Introduction
Tyukin (2015) established the first result for quenched coalescents, showing that in a well-mixed population in the absence of uneven offspring distributions, gene genealogies of unlinked loci really are independent in the large population limit. However, the works Diamantidis et al. (2024); Alberti et al. (2025) then showed the effects of uneven offspring distributions in the absence of selfing on gene genealogies. The pedigree leaves a persistent imprint on the dynamics of these gene genealogies through prodigious progenitors, individuals whose genetic contributions are shared across the genome. The limiting quenched genealogy becomes time inhomogeneous, evolving with an ordinary Kingman-like background punctuated by rare generations in which prodigious individuals contribute a non-vanishing proportion of the genetic material of the population even in the infinite population limit. Under the same assumptions of Birkner et al. (2018) that yield an annealed multiple merger coalescent, gene genealogies conditional on the pedigree converge to a coalescent process driven by a Poisson point process recording the timing and structure of exceptional reproductive generations of the pedigree. Between exceptional generations, lineages merge only in binary mergers at a constant rate. At exceptional generations, many lineages may merge simultaneously, with the merger pattern determined by the ordered offspring frequencies. This identifies the precise sense in which widely separated loci are coupled; they share the same record of macrosopic pedigree events, these prodigious progenitors. This Poisson driven viewpoint is closely related to the limiting object in the present paper.
@@
Much like the well-mixed case, the structured coalescent of Notohara (1990); Wilkinson-Herbots (1998) and multi-type $\Xi$ coalescents Eldon (2009); Möhle (2024); Pra et al. (2025) describe an averaged single locus genealogy. They do not condition on the single realized history of the pedigree. In structured populations this distinction becomes especially natural. Large migration waves can occur on the same time scale as coalescence and can affect many lineages at once. Such events are recorded in the pedigree and are shared across loci. They therefore create an additional source of time inhomogeneity and cross locus dependence that is invisible to purely annealed descriptions. Wilton et al. (2017) provided simulation results that demonstrated the importance of the pedigree in structured populations, where deviations from standard coalescent predictions can persist longer than in well-mixed populations due to the specific history of migration events captured by the pedigree.
@@ 1.1 The contributions of this paper
We control these joint laws by constructing a coupling of the $l$ conditionally independent copies on the same pedigree, in the same spirit as Alberti et al. (2025), who built on earlier work of Bolthausen and Sznitman (2002); Birkner et al. (2012). The coupling makes transparent the two mechanisms that drive the limit. Firstly, genuinely large pedigree events that change the coalescent state, such as macroscopic migration waves or unusually large families, are rare on the coalescent time scale. However, they affect all copies in nearly the same way. In particular, the different copies jump to corresponding new states with asymptotically matching probabilities. Second, during the overwhelming majority of small generations, the copies experience the “neutral” migration patterns described above, and their jump decisions become asymptotically independent across copies. This approximate independence follows from straightforward moment bounds. These two observations form the basis for a “separation of scales” argument in which we bridge these two regimes to demonstrate the desired convergence of the joint law. We note here that there is an implicit separation of scales argument in the $\varepsilon$-naive-coalescent of Alberti et al. (2025), though the total form of the argument is distinct. A strength of this method is that it only relies on some crude moment bounds and law of large numbers type arguments, which makes the analysis particularly simple and general.
@@ 2.2 The pedigree and its ancestral process
A partition with type in $F$, where $F$ is an arbitrary topological space, is a pair $(\xi,f)$, where $\xi\in\mathcal{E}_{n}$ and
@@
  • the blocks containing $i$ and $j$ are paired in $\xi$ if and only if $\widehat{X}_{i}^{N}(k)=\widehat{X}_{j}^{N}(k)$ and $M_{i}^{N}(k)\neq M_{j}^{N}(k)$,
@@ 3 An annealed scaling limit
for every $(v,w)$ in $E$. In particular, when $m_{(v,w)}$ is supported on $\{0,\frac{1}{N},\ldots,\frac{N-1}{N},1\}$ it is equivalent to
$\lim_{N\to\infty}\frac{1}{c_{N}^{v_{0}}}\mathbb{E}\left[m_{(v,w)}\right]=\bar{\mu}_{(v,w)}.$
@@
This intensity is such that any block begining in deme $v$ is thrown uniformly at random on the interval for deme $w$ with probability $m_{(v,w)}$ for $w\neq v$, and probability $1-\sum_{w\neq v}m_{(v,w)}$ for being thrown uniformly at random on the interval for deme $v$. Blocks are identified if their points lie in the same deme and in the same interval determined by $x^{v}$. Denote by $q_{n}(x,m)(\xi,\eta)$ the probability that an element $\xi$ coalesces into the element $\eta$, conditional on $(x,m)$.
@@ 4 Main results
The proof essentially proceeds by the method of moments. This is possible, due to a corollary of Fan et al. (2025), which guarantees that quenched convergence in finite-dimensional distribution of càdlàg paths is enough to guarantee weak convergence in distribution of quenched laws when the space the paths take values in is Polish and locally compact. We make this reasoning precise in the following lemma.
@@
$\mathbb{E}\left[\mathbb{P}\left(cd_{V}(\overline{\chi}^{N,n})\in C(\vec{t},\vec{\xi})\,\mid\,\mathcal{A}_{N}\right)^{l}\right]=\mathbb{P}\left(\left(cd_{V}\left(\overline{\chi}_{i}^{N,n}\right)\right)_{i=1}^{l}\in C(\vec{t},\vec{\xi})^{l}\right)$
@@
At the same time, when we restrict ourselves to the rare jumps where $(\widetilde{\mathcal{V}}(k),m(k))\in B(\varepsilon)^{c}$ we have that the transitions of the conditionally independent realizations are really asymptotically identical.
@@
Suppose, as $N\to\infty$, that Assumptions 3.2, 3.3, 3.4, 3.7, and 3.12 hold. Let
$P_{N,l}(t):=\left(\mathbb{P}\left(\left(cd_{V}\left(\overline{\chi}_{i}^{N,n}\right)\right)_{i=1}^{l}=\vec{\eta}\,\mid\,\left(cd_{V}\left(\overline{\chi}_{i}^{N,n}\right)(0)\right)_{i=1}^{l}=\vec{\xi}\right)\right)_{\vec{\xi},\vec{\eta}\in\mathcal{E}_{n}(V)^{l}}$
denote the transition kernel of $\left(cd_{V}\left(\overline{\chi}_{i}^{N,n}\right)\right)_{i=1}^{l}$ at time $t$. Then, for any fixed $T>0$, there is a continuous function $g_{T}:\mathbb{R}_{+}\to\mathbb{R}_{+}$ with $g(0)=0$ such that
@@
Finally, we characterize the annealed joint law of $l$ realizations $\left(\chi_{i}^{n}\right)_{i=1}^{l}$ of a $\Psi$-driven $n$-$(\kappa,\mu)$-coalescent $\chi^{n}$ that are conditionally independent with respect to $\Psi$. We show that $\left(\chi_{i}^{n}\right)_{i=1}^{l}$ has infinitesimal generator $\mathcal{L}$.
@@ 6.1 A two-deme Wright-Fisher model with neutral migration
$m_{(1,2)}=m_{(2,1)}\sim\frac{1}{N}\text{\emph{Binomial}}\left(N;\frac{1}{N}\frac{\sigma}{2}\right)$
@@ 6.2 A two-deme modified Wright-Fisher model with large offspring numbers
$m_{(1,2)}=m_{(2,1)}\sim\frac{1}{N}\text{\emph{Binomial}}\left(N;\frac{\sigma}{2}\frac{1}{N_{\gamma}}\right)$
@@
$\kappa_{\gamma}:=\begin{cases}\left(1,1\right),&\gamma>1,\\ \left(\dfrac{1}{1+\phi\psi^{2}},\dfrac{1}{1+\phi\psi^{2}}\right),&\gamma=1,\\ \left(0,0\right),&0<\gamma<1,\end{cases}\qquad\mu_{\gamma,(1,2)}=\mu_{\gamma,(2,1)}:=\begin{cases}\sigma,&\gamma>1,\\ \dfrac{\sigma}{1+\phi\psi^{2}},&\gamma=1,\\ 0,&0<\gamma<1.\end{cases}$
@@
We now identify the vague limit in Assumption 3.4. In a small generation, $\widetilde{\mathcal{V}}^{v}$ is Wright-Fisher reproduction, so its contribution vanishes on $\Delta^{V}\times[0,1]^{E}\setminus\{\mathbf{0}_{V,E}\}$ by (Birkner et al., 2018, Proposition 2.1). In a large generation in deme $v$, the $\lfloor\psi N\rfloor$ offspring of the prolific couple inherit, at a fixed locus, one of the two parental copies from each of the two parents independently; thus the four parental copies each contribute $\mathrm{Bin}\left(\lfloor\psi N\rfloor,1/2\right)$ gametes, and after normalisation by $2N$ this yields the limit $x_{\psi}$ in $\Delta$. Since large generations occur with probability $\varepsilon_{N}$ independently in the two demes, the event of simultaneous large generations has probability $\varepsilon_{N}^{2}=o(c_{N})$, and does not contribute in the $1/c_{N}$ scaling. Consequently,
@@
given by (6.3). This is exactly Assumption 3.4 with $\Phi=\Phi_{\gamma}$. By a direct computation one obtains $\mu_{\gamma,(1,2)}=\mu_{\gamma,(2,1)}$ as stated in the proposition and Assumption 3.7 holds.
@@ 6.3 A two-deme model with random individual fitness and large migrations
$\lim_{z\to\infty}\mathbb{P}\left(W\geq z\right)z^{\alpha}=c_{W}$
@@
$d\Lambda_{\alpha}(y)=\frac{1}{\Gamma(2-\alpha)\Gamma(\alpha)}\,y^{1-\alpha}(1-y)^{\alpha-1}\,dy,\qquad d\nu_{\alpha}(y):=\frac{d\Lambda_{\alpha}(y)}{y^{2}}.$
@@
$\Phi^{\mathrm{fit}}:=c_{\alpha}\int_{(0,1)}\left(\delta_{\left(x_{y}^{(1)},\mathbf{0}_{E}\right)}+\delta_{\left(x_{y}^{(2)},\mathbf{0}_{E}\right)}\right)\,d\nu_{\alpha}(y),$
@@
by (Birkner et al., 2018, Lemma 2.2). In particular, Assumptions 3.2 and 3.3 hold.